find four consecutive ODD integers such that four less than twice the second has the same value as three times the number that is one more than the third.
Answer by ankor@dixie-net.com(15628)

find four consecutive ODD integers
x, x+2, x+4, x+6
;
such that four less than twice the second has the same value as three times the number that is one more than the third.
2(x+2) - 4 = 3((x+4)+1)
2x + 4 - 4 = 3(x+5)
2x = 3x + 15
-15 = 3x - 2x
x = -15
:
The four numbers: -15, -13, -11, -9
:
:
Check solution in the given statement:
"four less than twice the second has the same value as three times the number that is one more than the third."
2(-13) - 4 = 3(-11 + 1)
-26 - 4 = 3(-10)